OpenStudy (anonymous):

Write the sum using summation notation, assuming the suggested pattern continues. -9 - 4 + 1 + 6 + ... + 66 Never learned what to do when missing a number...

thomaster (thomaster):

As you seem, 5 is added everytime

thomaster (thomaster):

-9 + 5 = -4 -4 + 5 = 1 1 + 5 = 6 So the next number will be 6+5 = 11